OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x)= 2/x and g(x)= 2/x

OpenStudy (anonymous):

Please please please help me!

OpenStudy (anonymous):

\[ g(f(x))=\frac {2}{f(x)}= \frac {2}{\frac 2 x}= 2 \frac {x}{2}=x\\ f(g(x))=\frac {2}{g(x)}= \frac {2}{\frac 2 x}= 2 \frac {x}{2}=x \]

OpenStudy (anonymous):

You're my savior :) Thanks!